4 equal charges of value Q are arranged at the four corners of a square of side a .A unit positive charge of mass M is placed at a height h above the centre of the square. for the equilibrium of the unit charge the value of q must be

Grade:12

1 Answers

Shriya Sanjeev Gupta
7 years ago
Given: 4 charges :q1=q2=q3=q4=QUnit positive charge=1C at center of square at height =hDistance of Q from center=a/√2. ( if side of square= a)So dist. Of Q from unit charge=√(a/√2)^2+h^2Resultant force from all 4 charges= 4Fcosø.therefore 4 Fcosø=mg=>(4×k×Q×1/(h^2+a^2/2))×h/√h^2+a^2/2=mg=> 4hkQ/(h^2+a^2/2)^3/2=mg=>Q=mg(h^2+a^2/2)^3/2/(4hk)SolvedRegards Shriya...